Anushka (an Australian Tax resident) works as an employee for a child care centre, Brilliant Kids Pty Ltd, on a permanent part-time basis while she runs her business as a day carer for her own customers. Meanwhile, she invested some of her extra cash in some shares on the ASX and in an Australian private company. Listed below is the summary of her earnings for the year ended 30 June 2020.
Particulars $
Net Salary Received from Brilliant Kids for the year (Tax paid by company to ATO during the year is $29,000)
81,000
Net Business Profit as a sole trader (Day Carer)
30,000
Allowable Deductions on Employment (Uniforms, Laundry, Shoes)
5,000
Fully Franked Dividend
7,000
Unfranked Dividend from a private company
1,200

Required:
Calculate Anushka’s taxable income and net tax payable (Ignore Medicare Levy and Medicare Levy Surcharge if any).
(Please use the following table to complete your response to this question in the spaces provided)
Particulars Answer Marks
Net Salary Received
0.5 mark
Add Back Tax Paid
0.5 mark
Net Business Profit
0.5 mark
Fully Franked Dividend
0.5 mark
Add Back Franking Credit
1 mark
Unfranked Dividend
0.5 mark
Assessable Income
1 mark
(Allowable Deductions)
0.5 mark
Taxable Income
1 mark
Tax Payable
2 marks
Less Refundable Tax Offsets
PAYG Withholding
0.5 mark
Franking Credits
0.5 mark
Net Tax Payable
1 mark
Total Marks
10 marks
Answer #1

Net salary received   81,000

Add back tax paid   29,000

Net business profit     30,000

Fully Franked dividend     7,000

Add back Franking credit   3,000

Unfranked dividend   1,200

Assessable income   151,200

(less) allowable deductions     5,000

Taxable income   146,200

Tax payable   41,507  

Less refundable tax offset PAYG  29000

Fracking credits   3,000

Net tax payable     $9,507

Workings

Tax payable- income lies on $90,001- $180,000.

Tax on unfranked dividend 30% of $1,200= $360

Balance taxable income 146200-1200 = 145,000

Tax calculation- $20,797+(37% x ($145,000 – $90,000)) = $41,147

Including unfranked dividend tax = 360+41147 = 41,507